NURBS 3-11-99 G. Farin

Programming assignment #3

due 3-25

Nonrational Bezier curves are not invariant under projective maps. Illustrate this concept using the circular model of the projective plane and the programs in the ftp site.

Hand in two different examples, each with illustration(s) and writeup.